### v1

# Fast Transaction Scheduling in Blockchain Sharding

Abstract.
Sharding is a promising technique for addressing the scalability issues of blockchain. It divides the $n$ participating nodes into $s$ disjoint groups called shards, where each shard processes transactions in parallel. The accounts (shared objects) are also distributed among the $s$ shards. We investigate scheduling algorithms for the blockchain sharding systems, where each transaction resides in a shard of the communication graph and attempts to access accounts at possibly remote shards. We examine batch scheduling problems on the shard graph $G_{s}$, where given a set of transactions, we aim to find efficient schedules to execute them as fast as possible. First, we present a centralized scheduler where one of the shards has global knowledge of transactions to be processed. For general graphs, where the transaction and its accessing objects are arbitrarily far from each other with a maximum distance $d$, the centralized scheduler provides $O(kd)$ approximation to the optimal schedule, where $k$ is the maximum number of shards each transaction accesses. Consequently, for a Clique graph where shards are at a unit distance from each other, we obtain $O(k)$ approximation to the optimal schedule. We also get $O(k\log s)$ approximation for Hypercube, Butterfly, and $g$-dimensional Grid, where $g=O(\log s)$. Next, we provide a centralized scheduler with a bucketing approach that offers improved bounds for special cases. For the General graph where each transaction chooses $k$ objects randomly we give $O(k\log D\cdot(k+\log s))$ approximation, and for the Line graph we provide $O(k\sqrt{d}\log D)$ approximation to the optimal schedule, where $D$ is the diameter of shard graph $G_{s}$. Finally, we provide a distributed scheduler where shards do not require global transaction information. We achieve this by using a hierarchical clustering of the shards and using the centralized scheduler in each cluster. We show that the distributed scheduler has a competitive ratio of $O(\mathcal{A_{\mathcal{CS}}}\log^{2}s)$, where $\mathcal{A_{\mathcal{CS}}}$ is the approximation ratio of the centralized scheduler. To our knowledge, we are the first to give provably fast transaction scheduling algorithms for blockchain sharding systems.
Blockchain, Blockchain Sharding, Transaction Scheduling, Batch Scheduling.
ccs: Computing methodologies Distributed algorithmsccs: Theory of computation Scheduling algorithms.
1. Introduction
Blockchains are known for their special features such as fault tolerance, transparency, non-repudiation, immutability, and security (Sankar et al., 2017). Thus, blockchains have been used in various domains, such as cryptocurrency (Nakamoto, 2009; Buterin et al., 2014), healthcare (McGhin et al., 2019), digital forensics (Troyer et al., 2021; Akbarfam et al., 2023b, a), supply chain management (Azzi et al., 2019), edge and scientific computing (Al-Mamun et al., 2022, 2021).
Blockchain technology is also gaining interest from banks and financial institutions because of its transformational power in the financial industry (Androulaki et al., 2020; Javaid et al., 2022; Alamsyah et al., 2024; Weerawarna et al., 2023). This innovation is expected to reshape financial services, revolutionize banking operations, and address dynamic industry demands without sacrificing privacy (Baum et al., 2023; Weerawarna et al., 2023). Implementing blockchain can offer competitive benefits to traditional banking systems by enhancing transaction privacy and security at a minimum cost (Cucari et al., 2022). Traditional financial institutions, tech firms, and startups are also coordinating with financial technology to offer affordable and easy to use financial services (Canaday, 2017). However, the primary drawback of blockchain is its scalability and its performance, which is dependent on the throughput of the transactions (Cocco et al., 2017; Chen and Bellavitis, 2020). Thus, there is a research gap to address the scalability issues of blockchain.
To append a new block in the blockchain network, each participating node needs to reach a consensus, which is a time and energy-consuming process (Adhikari and Busch, 2023; Adhikari et al., 2024a, b). Moreover, each node is required to process and store all transactions, which leads to scalability issues in the blockchain system. To improve scalability and performance, sharding protocols have been proposed, such as Elastico (Luu et al., 2016), OmniLedger (Kokoris-Kogias et al., 2018), RapidChain (Zamani et al., 2018), Sharper (Amiri et al., 2021), and ByShard (Hellings and Sadoghi, 2021). However, these sharding protocols do not provide a formal analysis of how we can schedule the transaction as fast as possible. Sharding divides participating nodes into smaller groups called shards that allow the parallel processing of transactions. In the sharded blockchain, independent transactions are processed and committed in multiple shards concurrently, which improves the blockchain system’s throughput.
We consider a blockchain system consisting of $n$ nodes, which are further divided into $s$ shards, and each shard consists of a set of nodes. Shards are connected in a graph network $G_{s}$ with a diameter $D$, and each shard holds a subset of the objects (accounts). We assume that transactions are distributed across the shards, and each transaction accesses at most $k$ accounts. A transaction $T$ initially is in one of the shards; we called it as home shard for $T$. Similar to other sharding systems (Hellings and Sadoghi, 2021; Adhikari and Busch, 2023; Adhikari et al., 2024a, b), each transaction $T$ is split into subtransactions, where each subtransaction accesses an account. A subtransaction of $T$ is sent to the destination shard that holds the respective account. The maximum distance between the home shard and the destination shards in the shard graph is at most $d\leq D$. In other words, the transaction and its accessing shards (objects) are arbitrarily far from each other with a maximum distance of $d$. All home shards process transactions concurrently. A problem occurs when conflicting transactions try to access the same account simultaneously. In such a case, the conflict prohibits the transactions to be committed concurrently and forces them to serialize (Adhikari et al., 2024a, b). Our proposed scheduling algorithms coordinate the home shards and destination shards to process the transactions (and respective subtransactions) in a conflict-free manner. Each destination shard maintains a local blockchain of the subtransactions that are sent to it. The global blockchain can be constructed by taking the union of the local blockchains at the shards (Adhikari and Busch, 2023).
We consider batch problem instances where a (batch) set of transactions ${\mathcal{T}}$ initially reside at shards and need to be processed. We study the shard model in different graph topologies such as General graph, Clique (Complete) graph, Hypercube, Butterfly, Grid, and Line graphs and provide the analysis for execution time. The scheduler determines the time step for each transaction $T_{i}\in{\mathcal{T}}$ to process and commit. The execution of our scheduling algorithm is synchronous, where the timeline is divided into time steps referred to as rounds. Similar to the work in (Adhikari et al., 2024a, b), we consider that the duration of a round is sufficient to allow the execution of the PBFT consensus algorithm (Castro et al., 1999) in each shard. A round is also the time to send a message between shards in a unit distance. The main goal of a scheduling algorithm is to efficiently and fairly process all transactions while minimizing the total execution time (makespan) until all transactions are either committed or aborted. Unlike previous sharding approaches (Hellings and Sadoghi, 2021; Kokoris-Kogias et al., 2018; Zamani et al., 2018), our scheduling algorithms do not require locking mechanisms for concurrency control. Hence, we can also say our algorithms are lock-free.
Shards Connected as Approximation to Optimal Schedule of
Centralized Scheduler Distributed Scheduler
General Graph $O(kd)$ $O(kd\log^{2}s)$
Hypercube, Butterfly, and
$g$-dimensional Grid Graph $O(k\log s)$ $O(k\log^{3}s)$
General Graph where,
$k$ is chosen as random $O(k\log D\cdot(k+\log s))$ $O(k\log D\cdot(k+\log s)\log^{2}s)$
Line Graph $O(k\sqrt{d}\log D)$ $O(k\sqrt{d}\log D\cdot\log^{2}s)$
Table 1. Summary of the proposed centralized and distributed scheduler approximation to the optimal schedule. The used notations are as follows: $s$ represents the total number of shards, $k$ denotes the maximum number of shards (objects) accessed by each transaction, $d$ denotes the worst distance between any transaction (home shard) and its accessing objects (destination shard), $D$ denotes the diameter of shard graph $G_{s}$.
1.1. Contributions and Techniques
To our knowledge, we give the first provably efficient batch transaction scheduling algorithms for blockchain sharding systems. We provide a summary of contributions in Table 1, which are also described as follows:
  • Centralized Scheduler: First, we provide a Centralized scheduler where one of the shards has global knowledge of the transactions. Consider a general graph $G_{s}$ with diameter $D$. Suppose that each transaction accesses at most $k$ objects, where the distance between a transaction and its accessing objects is at most $d\leq D$. The centralized scheduler provides an $O(kd)$ approximation to the optimal schedule. Hence, when $G_{s}$ is a Clique graph, where shards are at a unit distance from each other (thus, $d=1$), we obtain an $O(k)$ approximation to the optimal schedule. Moreover, we obtain an $O(k\log s)$ approximation for the Hypercube, Butterfly, and $g$-dimensional Grid where $g=O(\log s)$, since in all these cases $d=O(\log s)$.
    • Technique: Our scheduling algorithm consists of two main phases. In the first phase, the scheduler determines the order of execution of the transactions by coloring transactions that need to be processed using a conflict graph. In the second phase, our algorithm confirms and commits each transaction according to the schedule determined by the first phase. In the centralized scheduler, we consider a set of transactions ${\mathcal{T}}$ that are distributed in a shard graph $G_{s}$, and one of the shard as leader shard are given to the algorithm as input. Then, the scheduler provides the schedule and processes transactions ${\mathcal{T}}$. We assume that the leader shard has global knowledge of the transactions that need to proceed. Suppose $G_{\mathcal{T}}$ is the transactions conflict graph where each node represents a transaction, and an edge between two nodes corresponds to a conflict between them. In the first phase, the leader shard schedules the transactions using a greedy coloring of $G_{\mathcal{T}}$, which assigns execution times to each transaction based on the coloring of $G_{\mathcal{T}}$. A valid coloring of $G_{\mathcal{T}}$ assigns a unique positive integer to each transaction such that two adjacent transactions (conflicting transactions) receive different colors. The colors correspond to the distinct time steps (we also called round) where the respective transactions are processed for commit. After coloring, the leader shard shares the color information with each home shard. Then, in the second phase, each home shard processes its transactions to commit in the destination shards. We described the proposed centralized scheduling algorithm in Section 4.
  • Centralized Scheduler with Buckets: Next, we give the centralized scheduler with a bucketing approach. Similar to a centralized scheduler, we consider a shard graph $G_{s}$ with diameter $D$, and the maximum distance between the home shard of the transaction and its accessing objects is at most $d\leq D$. We divide the transactions into buckets $B_{0},B_{1},\ldots$, according to the maximum distance between the home shard of the transaction and its accessing objects (i.e., destination shards). $B_{i}$ consists of the transactions whose maximum distance between the home shard and its destination shard is in range $[2^{i},2^{i+1})$. This approach offers improved bounds for special cases: for the General graph where the $k$ objects are chosen randomly per transaction we give $O(k\log D\cdot(k+\log s))$ approximation, and for the Line graph, we provide $O(k\sqrt{d}\log D)$ approximation to the optimal schedule.
    • Technique: Each home shard of transaction $T_{i}\in{\mathcal{T}}$ discovers the position of its accessing objects (destination shards) and sends that transaction information to the respective bucket $B_{i}$. Bucket $B_{i}$ denotes the set of transactions whose distance between the home shard and the accessing objects (destination shards) is in the range $[2^{i},2^{i+1})$; $i\in\{0,\dots,\lfloor\log D\rfloor\}$ where $D$ is the diameter of shards graph $G_{s}$. We assume that each bucket $B_{i}$ has its own leader shard $\ell_{i}$, which is known to all shards in $G_{s}$. First, each transaction determines its bucket, and then we run the centralized scheduler for each bucket. Details of the algorithm is provided in Section 5.
  • Distributed Scheduler: Finally, we give a distributed scheduler where shards are not aware of global transaction information. We use a clustering approach for the shard network by hierarchical clustering (Gupta et al., 2006) so that independent transactions can be scheduled and committed concurrently. The basic idea is to simulate the centralized scheduler within each cluster of the hierarchy to make a distributed scheduler. We show that the distributed scheduler has a competitive ratio of $O(\mathcal{A_{\mathcal{CS}}}\log^{2}s)$, where $\mathcal{A_{\mathcal{CS}}}$ is the approximation ratio of the centralized scheduler.
    • Technique: We consider that our distributed scheduling algorithm uses a hierarchical decomposition of shard graph $G_{s}$ which is known to all the shards and calculated before the algorithm starts. This shard clustering (graph decomposition) is based on the clustering techniques in (Gupta et al., 2006) and which were later used in (Sharma and Busch, 2014; Busch et al., 2022; Adhikari et al., 2024a, b). We divide the shard graph $G_{s}$ into the hierarchy of clusters with layers and sublayers. The home cluster for each transaction $T_{i}$ is defined as follows: suppose $S_{i}$ is the home shard of $T_{i}$, and $z$ is the maximum distance from $S_{i}$ to the destination shards that will be accessed by $T_{i}$; the home cluster of $T_{i}$ is the lowest-layer (and lowest sub-layer) cluster in the hierarchy that contains the $z$-neighborhood of $S_{i}$. Each home cluster has one dedicated leader shard, which will handle all the transactions that have their home shard in that cluster (i.e., transaction information is sent from the home shard to the cluster leader shard for processing). After that, each cluster leader shard schedules the transaction of its cluster by assigning color using a greedy coloring algorithm (similar to the centralized scheduler). To achieve the benefit of locality and reduce the transaction latency, the lower cluster layer gets priority in processing its transactions. Within each cluster, we simulate and run a centralized scheduler algorithm to schedule transactions. A detailed description of our distributed scheduler is in Section 6.
Paper Organization: The rest of this paper is structured as follows: Section 2 provides related works. Section 3 describes the preliminaries for this study and the sharding model. Section 4 presents a centralized scheduling algorithm. In Section 5, we provide a centralized scheduler with the bucketing approach. Section 6 generalizes the techniques to a fully distributed setting. Finally, we give our conclusions in Section 7.
2. Related Work
Extensive research has been proposed to address the scalability issues of blockchain in the consensus layer (Jalalzai and Busch, 2018; Jalalzai et al., 2019, 2021). While these proposed algorithms and protocols have made some improvements in scalability, the system’s performance still degrades as the network size increases. To tackle the scalability issue of blockchain, various sharding protocols (Luu et al., 2016; Kokoris-Kogias et al., 2018; Zamani et al., 2018; Hellings and Sadoghi, 2021; Dang et al., 2019; Amiri et al., 2021; Adhikari and Busch, 2023) have been proposed. These protocols have shown promising enhancements in the transaction throughput of blockchain by processing non-conflicting transactions in parallel in multiple shards. However, none of these protocols have specifically explored the batch scheduling of transactions within a sharding environment, and they do not provide an extensive theoretical analysis for transaction scheduling.
To process transactions parallelly in the sharding model, some research work (Kokoris-Kogias et al., 2018; Hellings and Sadoghi, 2021; Set and Park, 2022) has used two-phase locking for concurrency control. However, locks are expensive because when one process locks shared data for reading/writing, all other processes attempting to access the same data set are blocked until the lock is released, which lowers system throughput. Our scheduling algorithm does not use locks, as concurrency control is managed by the scheduler by processing non-conflicting transactions concurrently. In (Adhikari and Busch, 2023) the authors propose lockless blockchain sharding using multi-version concurrency control. However, they lack an analysis and technique for optimal scheduling. Moreover, they do not provide the benefits of locality and optimization techniques for transaction scheduling.
Several works have been conducted on transaction scheduling in shared memory multi-core systems, distributed systems, and transactional memory. In a recent work (Busch et al., 2023), a stable scheduling algorithm was proposed specifically for software transactional memory systems, considering transactions generated by an adversarial model. Similarly, in research works (Attiya et al., 2015; Sharma and Busch, 2014, 2015), authors explored transaction scheduling in distributed transactional memory systems aimed at achieving better performance bounds with low communication costs. Authors in work (Busch et al., 2017) provided offline scheduling for transactional memory, where each transaction attempts to access an object, and once it obtains the object, it executes the transaction. In another (Busch et al., 2022), the author extended their work from offline to online scheduling for the transactional memory model. However, these works do not address transaction scheduling problems in the context of blockchain sharding. This is because, in the transactional memory model, the object is mobile, and once the transaction obtains the object, it immediately executes the transaction. In contrast, in blockchain sharding, the object is static in the shard, and there is a confirmation scheme to confirm and commit each subtransaction consistently in the respective accessing shard.
In a recent work (Adhikari et al., 2024a, b), the authors provide a stability analysis of blockchain sharding considering adversarial transaction generation. However, their solution does not consider different types of sharding graphs, and they also do not provide a theoretical analysis of the optimal approximation for the scheduling algorithm.
3. Technical Preliminaries
A block is a data structure that consists of a set of transactions. The block header contains additional metadata, including the block hash, previous block hash, block sequence, etc. Blockchain is simply a chain of blocks, where a block points to its predecessor (parent) through its hash, which makes them immutable. A blockchain is essentially implemented as a decentralized peer-to-peer ledger where the blockchain is replicated across multiple interconnected nodes.
3.1. Blockchain Sharding Model
We consider blockchain sharding model which is similar to  (Adhikari and Busch, 2023; Hellings and Sadoghi, 2021; Adhikari et al., 2024a, b), consisting of $n$ nodes which are partitioned into $s$ shards $S_{1},S_{2},\dots,S_{s}$ such that $S_{i}\subseteq\{1,\ldots,n\}$, for $i\neq j$, $S_{i}\cap S_{j}=\emptyset$, $n=\sum_{i}|S_{i}|$, and $n_{i}=|S_{i}|$ denotes the number of nodes in shard $S_{i}$. Let $G_{s}=(V,E,w)$ denote a weighed graphs of shards, where $V=\{S_{1},S_{2},\dots,S_{s}\}$, and the edges $E$ correspond to the connections between the shards, and the weight represents the distance between the shards.
Shards communicate with each other via message passing, and here, we are not focusing on optimizing the message size. Moreover, all honest nodes in a shard agree on each message before transmission (e.g. running the PBFT (Castro et al., 1999) consensus algorithm within the shard). Similar to work in (Hellings and Sadoghi, 2021, 2022; Adhikari et al., 2024a, b), we assume that we are given a cluster-sending protocol for reliable and secure communication between shards, satisfying the following properties for sending data $\mathfrak{D}$ from shard $S_{x}$ to shard $S_{y}$:
  • Shard $S_{x}$ sends $\mathfrak{D}$ to $S_{y}$ if there is an agreement among the honest nodes in $S_{x}$ to send $\mathfrak{D}$;
  • All honest nodes in recipient shard $S_{y}$ will receive the same data $\mathfrak{D}$;
  • All non-faulty nodes in sender shard $S_{x}$ receive confirmation of data $\mathfrak{D}$ receipt.
We assume that the above cluster-sending properties between two shards at unit distance are guaranteed to be satisfied within a single communication round. We assume that the communication is synchronous and time is divided into rounds. Moreover, the distance between two shards is measured as the number of rounds required to send a message from one shard to another over the graph $G_{s}$. Note that some of the messages need to propagate through multiple shards in $G_{s}$.
Similar to work in (Adhikari et al., 2024a, b), for secure inter-shard communication, we use a broadcast-based protocol that can operate in a system with Byzantine failures (Hellings and Sadoghi, 2022). For the communication between shard $S_{1}$ and $S_{2}$, shard $S_{1}$ uses a consensus protocol to reach agreement on a value $v$. Then, a set $M_{1}\subseteq S_{1}$ of $f_{1}+1$ nodes in $S_{1}$ and a set $M_{2}\subseteq S_{2}$ of $f_{2}+1$ nodes in $S_{2}$ are chosen, where $f_{i}$ is the number of faulty nodes in shard $S_{i}$. Finally, each node in $M_{1}$ is instructed to broadcast $v$ to all nodes in $M_{2}$. Due to the size of $M_{1}$ and $M_{2}$, at least one non-faulty node in $S_{1}$ will send a value to a non-faulty node in $S_{2}$, which is sufficient to receipt and confirmation of $v$ in $S_{2}$.
Each shard consists of a local blockchain (which is part of the global blockchain) according to the subtransactions it receives and commits. We use $f_{i}$ to represent the number of Byzantine nodes in shard $S_{i}$. To guarantee consensus on the current state of the local blockchain, we assume that every shard executes the PBFT (Castro et al., 1999) consensus algorithm. In order to achieve Byzantine fault tolerance, each shard $S_{i}$ consists of $n_{i}>3f_{i}$ nodes, where $n_{i}$ is the total number of nodes in $S_{i}$.
Suppose we have a set of shared accounts $\mathcal{O}$ (which we also call objects). Similar to previous work in (Hellings and Sadoghi, 2021; Adhikari and Busch, 2023; Adhikari et al., 2024a), we assume that each shard is responsible for a specific subset of the shared objects (accounts). To be more specific, $\mathcal{O}$ is split into disjoint subsets $\mathcal{O}_{1},\ldots,\mathcal{O}_{s}$, where the set of items under the control of shard $S_{i}$ is represented by $\mathcal{O}_{i}$. Every shard $S_{i}$ keeps track of local subtransactions that use its corresponding $\mathcal{O}_{i}$ to access objects.
We assume that shards are connected in a communication graph, and the diameter of the graph is $D$. Moreover, we assume that the distance between the transaction and its accessing objects are in ranges from $1$ to $d\leq D$.
3.2. Transactions and Subtransactions
Let’s consider a set of transactions ${\mathcal{T}}=\{T_{1},T_{2},\ldots\}$ that are distributed across different shards. Let’s say a transaction $T_{i}$ resides in a node $v_{T_{i}}$ within the system. The home shard of $T_{i}$ is the shard containing $v_{T_{i}}$. In this work, we consider the batch transactions that are already in the shard and need to proceed.
Similar to work in (Hellings and Sadoghi, 2021; Adhikari and Busch, 2023; Adhikari et al., 2024a, b), we define a transaction $T_{i}$ as a group of subtransactions $T_{i,a_{1}},\ldots,T_{i,a_{j}}$. Each subtransaction $T_{i,a_{l}}$ accesses objects only in $\mathcal{O}{a_{l}}$ and is associated with shard $S{a_{l}}$. Therefore, each subtransaction $T_{i,a_{l}}$ has a respective destination shard $S_{a_{l}}$. The home shard of $T_{i}$ sends subtransaction $T_{i,a_{l}}$ to shard $S_{a_{l}}$ for processing, where it is appended to the local blockchain of $S_{a_{l}}$. The subtransactions within a transaction $T_{i}$ are independent, meaning they do not conflict and can be processed concurrently. Like previous work in (Hellings and Sadoghi, 2021; Adhikari et al., 2024a, b), each subtransaction $T_{i,a_{l}}$ consists of two parts: (i) a condition check, where it verifies whether a condition of the objects in $\mathcal{O}{a_{l}}$ is satisfied, and (ii) the main action, where it updates the values of the objects in $\mathcal{O}{a_{l}}$.
Example 0.
Consider a transaction $T_{1}$ that involves read-write operations on accounts with certain conditions. For instance, let $T_{1}$ be “Transfer 3000 from Ray’s account to Alex’s account if Ray has 8000, Alex has 500 and Balen has 700”. The home shard of $T_{1}$ divides this transaction into three subtransactions $T_{1,r},T_{1,a},T_{1,b}$, where the destination shards $S_{r}$, $S_{a}$, and $S_{b}$ handle Ray’s, Alex’s, and Balen’s accounts respectively:
  • $T_{1,r}$
    - condition: “Check if Ray has 8000”
  • - action: “Deduct 3000 from Ray’s account”
  • $T_{1,a}$
    - condition: “Check if Alex has 500”
  • - action: “Add 3000 to Alex’s account”
  • $T_{1,b}$
    - condition: “Check if Balen has 700”
The home shard of $T_{1}$ sends the subtransactions to their respective destination shards. If the conditions are met (for example, in $T_{1,b}$ if Balen has 700) and the transactions are valid (for example, in $T_{1,r}$ if Ray indeed has 3000 in the account to be deducted), then the destination shards are prepared to commit the subtransactions in the local blockchains, which implies the entire transaction $T_{1}$ implicitly commits as well. However, if any of the conditions in the subtransactions are not met or the subtransactions are invalid, then all the subtransactions must abort (i.e., they are not added to the local blockchains), resulting in the abortion of $T_{1}$ as well.
Definition 0 (Conflict).
Transactions $T_{i}$ and $T_{j}$ are said to be conflict if they access the same object $o_{z}\in\mathcal{O}$ and at least one of these transactions writes (updates) the value of object $o_{z}$.
Transactions that conflict should be processed in a sequential manner to guarantee atomic object update. In such a case, their respective subtransactions should serialize in the exact same order in every involved shard. To resolve the conflict between two transactions $T_{i}$ and $T_{j}$ while accessing an object $o_{z}$, we schedule them one after another in such a way that $T_{i}$ executes before $T_{j}$ or vice versa.
4. Centralized Scheduler
We consider a shard graph $G_{s}$ built from $s$ shards, where a set of transactions ${\mathcal{T}}$ and a set of objects ${\mathcal{O}}$ are distributed across the $G_{s}$. In this centralized scheduler, we consider one shard as the leader shard $\ell$, which has global knowledge of the transactions that need to be processed. The leader shard runs a greedy coloring algorithm to schedule transactions in a conflict-free manner. The parameters ($G_{s},{\mathcal{T}},\ell$) are given as input to the algorithm, which then produces a schedule and processes the set of transactions ${\mathcal{T}}$. The details appear in Algorithm 1.
Consider each transaction $T_{i}\in{\mathcal{T}}$ accesses $k$ objects from a set of objects ${\mathcal{O}}$. Suppose $G_{\mathcal{T}}$ is the transactions interference (conflict) graph where each node represents a transaction and an edge between two nodes correspond to a conflict between them. The conflict occurs when the respective transactions try to access the same account (object) $o_{j}\in{\mathcal{O}}$ in the same shard. In Phase 1 of Algorithm 1, our scheduler schedules these transactions using a greedy vertex coloring of $G_{{\mathcal{T}}}$, which assigns execution times to each transaction based on the coloring of $G_{\mathcal{T}}$. A valid coloring of $G_{\mathcal{T}}$ assigns a unique positive integer to each transaction such that two adjacent transactions (conflicting transactions) receive different colors. The colors correspond to the distinct time steps (also called rounds) where the respective transactions are processed for commit. For each color, our algorithm takes four steps to confirm and commit the transaction in each destination shard, which is also shown in Phase $2$ of Algorithm 1. We continue to describe each phase of the algorithm as follows:
1 // txn: transaction; txns: transactions
2
Input : Shard graph $G_{s}$; set of txns ${\mathcal{T}}$ that need to be processed in $G_{s}$; $\ell$ is the leader shard for txns ${\mathcal{T}}$;
3
4
5// Phase 1: Graph Coloring and determining schedule
6
7Home shard of each txn $T_{i}\in{\mathcal{T}}$ shares the txn information with the leader shard $\ell$;
8
9Leader shard $\ell$ constructs $G_{\mathcal{T}}$, and runs a greedy coloring on $G_{\mathcal{T}}$ which returns $\xi$ colors;
10Leader shard $\ell$ informs the home shards of each txn $T_{i}\in{\mathcal{T}}$ about the respective txn colors;
11
12// Phase 2: confirming and committing
13 /* processed colored txn at appropriate round(s) */
14 for color $clr\leftarrow 1$ to $\xi$ do
15       /* Step 1: Home Shard */
16       The home shard of txn $T_{i}\in{\mathcal{T}}$ of color $clr$, splits $T_{i}$ into subtransactions ($T_{i,j}$) and sends them to respective destination shards for voting;
17
18      /* Step 2: Destination shards */
19       At destination shard $S_{j}$, if $T_{i,j}$ is valid and its condition is satisfied, then $S_{j}$ sends commit vote for $T_{i,j}$ to home shard of $T_{i}$; otherwise, $S_{j}$ sends abort vote for $T_{i,j}$ to home shard of $T_{i}$;
20
21      /* Step 3: Home Shard */
22       The home shard of $T_{i}$ checks whether all commit votes are received for $T_{i}$ from its destination shards; if true, the home shard sends confirm commit to destination shards; otherwise (if any abort vote received), the home shard sends confirm abort to destination shards;
23
24      /* Step 4: Destination Shards */
25       At destination shard $S_{j}$ of $T_{i,j}$, if $S_{j}$ receives confirm commit for $T_{i,j}$ (from home shard of $T_{i}$), then $S_{j}$ commits $T_{i,j}$ by appending it in the local blockchain; Otherwise, if $S_{j}$ receives confirmed abort (from home shard of $T_{i}$), then $S_{j}$ aborts $T_{i,j}$;
26
Algorithm 1 CentralizedScheduler ($G_{s}$, ${\mathcal{T}}$, $\ell$)
Phase 1: In this phase, each home shard of transaction $T_{i}\in{\mathcal{T}}$ sends its transaction information to the leader shard $\ell$. Then, the leader shard constructs the transactions conflict graph $G_{\mathcal{T}}$ and runs the greedy coloring algorithm on $G_{\mathcal{T}}$. The greedy coloring algorithm assigns the colors to each transaction of $G_{\mathcal{T}}$ using at most $\xi$ colors. After that, the leader shard informs the home shard of each transaction $T_{i}\in{\mathcal{T}}$ about the respective transaction color.
Phase 2: In this phase, each transaction is processed based on a computed schedule using the greedy coloring in Phase 1. Each home shard knows its transaction committing time and proceeds to determine whether it can be committed or not. Each color in $\xi$ takes $4$ steps for the confirmation (i.e., $4$ communication interactions between home shard and destination shards). So, for each color $clr\in\xi$ in Step 1, each home shard of transaction $T_{i}\in{\mathcal{T}}$ of color $clr$ splits the transaction $T_{i}$ into subtransactions ($T_{i,j}$) and sends them to the respective destination shards parallelly to check whether that transaction (i.e. its subtransactions) can be committed or not. In Step 2, each destination shard $S_{j}$ checks the conditions and validity of the subtransactions $T_{i,j}$ (such as account balance) and sends a commit vote to the home shard of $T_{i}$ if the subtransaction $T_{i,j}$ is valid; otherwise sends an abort vote to home shard of $T_{i}$ if the subtransaction $T_{i,j}$ is invalid. In Step 3, the home shard of $T_{i}$ collects the votes, and if it receives all the commit votes for a transaction $T_{i}$ from its destination shards, then it sends the confirmed commit message to the respective destination shards. Otherwise, if the home shard of $T_{i}$ receives any abort vote, then it will send a confirmed abort message to all respective destination shards. In Step 4, each destination shard either commits the subtransactions $T_{i,j}$ and appends them to a local ledger or aborts the subtransactions $T_{i,j}$ according to the message received from the home shard of $T_{i}$.
4.1. Analysis for General Graph
Suppose we run Algorithm 1 for the general graph whose diameter is $D$ and the distance between the home shard (transaction) and it’s accessing objects (destination shards) ranges from $1$ to $d$, where $d\leq D$. In the following section, we provide an analysis of such a case.
Lemma 0 (Lower Bound).
For the general graph, the processing time of any schedule for ${\mathcal{T}}$ in a graph $G_{\mathcal{T}}$ is at least $\Omega(l)$, where $l$ is the maximum number of transactions that access any object $o_{j}\in{\mathcal{O}}$.
Proof.
Let ${\mathcal{T}}_{o_{j}}$ denote the set of transactions that use object $o_{j}\in{\mathcal{O}}$. Let $l_{o_{j}}=|{\mathcal{T}}_{o_{j}}|$, and $l=\max_{j}l_{o_{j}}$ be the maximum degree in the graph for any object in ${\mathcal{O}}$. Then, the length of the schedule is at least $l$ because an object has to be accessed once at a time by $l$ transactions. Consequently, the processing time of any schedule for ${\mathcal{T}}$ is at least $\Omega(l)$. ∎
Lemma 0 (Upper Bound).
For the general graph, the time to process the transactions ${\mathcal{T}}$ is at most $O(kld)$, where $l$ is the maximum number of transactions that access an object $o_{j}\in{\mathcal{O}}$, $k$ is the maximum number of objects accessed by each transaction, and $d$ is the maximum distance between a transaction and its accessing objects.
Proof.
The maximum degree in the conflict graph $G_{\mathcal{T}}$ is at most $kl$, since each transaction accesses $k$ objects and each object is accessed by at most $l$ transactions. Therefore, the conflict graph $G_{\mathcal{T}}$ can be greedily colored with at most $\xi\leq kl+1$ colors.
The distance between a transaction and its accessing object is at most $d$ far away. Since Algorithm 1 has at most $4$ rounds of interactions between a transaction and its accessing objects in Phase 2, each color corresponds to $4d$ communication rounds. Thus, it takes at most $(kl+1)4d=O(kld)$ rounds to confirm and commit the transactions.
Theorem 3.
For the general graph, Algorithm 1 gives an $O(kd)$ approximation to the optimal schedule.
Proof.
From Lemma 2, the time to process all the transactions is at most $O(kld)$. Moreover, from Lemma 1, $\Omega(l)$ is the lower bound, which gives us $O(kd)$ approximation to the optimal schedule.
4.2. Analysis for Specialized Graphs
Let us consider shards are connected in a clique (complete) graph $G_{s}$ with $s$ shards where every shard is connected to every other shard with a unit distance (thus, $d=1$). Then, from Theorem 3, Algorithm 1 gives $O(k)$ approximation to the optimal schedule.
Let $G_{s}$ be a hypercube graph (Leighton, 2014). Then, there is a connecting path between each pair of shards with length $\log s$; hence $D=\log s$, and $d\leq D=\log s$. Therefore, Theorem 3 implies that there is an execution schedule which is a $O(k\log s)$ approximation of the optimal schedule.
Similar to the Hypergraph (Leighton, 2014), we can use the same result in the Butterfly networks (Leighton, 2014) and $g$-dimensional grids (Chan, 1989), where $g=O(\log s)$, which also gives an approximation for execution time as $O(k\log s)$.
Hence, from Theorem 3 we obtain the following corollaries.
Corollary 0.
For a clique graph with unit distance per edge (i.e. $d=1$), Algorithm 1 gives an $O(k)$ approximation to the optimal schedule.
Corollary 0.
Algorithm 1 gives an $O(k\log s)$ approximation for Hypercube, Butterfly, and $g$-dimensional grids graph, where $g=O(\log s)$.
5. Centralized Scheduler with Buckets
In the Centralized scheduler (Algorithm 1), a single shard (the leader shard) colors the transactions and determines the schedule. We propose an alternative algorithm that uses buckets to organize transactions based on the distances between the transaction and its accessed objects. This approach offers improved bounds for special cases when the $k$ objects are selected randomly, as well as for the Line graph.
Here, we distribute the coloring load in different shards, i.e., each bucket has its own leader shard, which is known to all the shards and determines the schedule to process the transaction of that bucket. We divide the transactions into buckets $B_{0},B_{1},\ldots$, according to the maximum distance between the home shard of the transaction and its accessing objects (i.e., destination shards). In bucket $B_{i}$, we add transactions whose maximum distance between home shard and its destination shard is in range $[2^{i},2^{i+1})$. We compute a schedule for each bucket. We also prioritize the processing of the lower index bucket transactions over the higher index bucket so that we can get the locality benefits and reduce transaction latency, that is, first, we calculate a schedule and process transactions of $B_{0}$, then for $B_{1}$, then for $B_{2}$, and so on.
The pseudocode of the centralized scheduler with the bucketing approach is presented in Algorithm 2. We assume that shard graph $G_{s}$, a set of transactions ${\mathcal{T}}$ that need to be processed, are given as input to the algorithm. Then, the scheduler provides the schedule and processes all the transactions ${\mathcal{T}}$. We assume that each bucket $B_{i}$ has its own leader shard $\ell_{i}$, which is known to all shards in $G_{s}$. First, each transaction determines its bucket, and then we run Algorithm 1 for each bucket by providing the shard graph $G_{s}$, set of transactions of each bucket $B_{i}$ and leader of that bucket as parameters.
1
Input : Shard graph $G_{s}$; set of txns ${\mathcal{T}}$;
2
[…]
4 Each Bucket $B_{i}$ has its own leader shard $\ell_{i}$ which is known to all shards in $G_{s}$;
5
6/* Run Algorithm 1 for each bucket $B_{i}$/
7 for Bucket $B_{i}\leftarrow B_{0}$ to Bucket $B_{\lfloor\log D\rfloor}$ do
8       Run CentralizedScheduler($G_{s}$, $B_{i}$, $\ell_{i}$);
9
10
Algorithm 2 SchedulerWithBuckets
5.1. Analysis for General Graph with Objects Chosen Randomly
Suppose each shard contains one transaction that needs to be processed and each transaction accesses at most $k$ objects picked randomly and uniformly from different shards. We use the following version of Chernoff bound in our analysis.
Lemma 0 (Chernoff bound).
Let $X_{1},X_{2},\cdots,X_{m}$ be independent Poisson trials such that, for $1\leq i\leq m$, $X_{i}\in\{0,1\}$. Then, for $X=\sum_{i=1}^{m}X_{i}$, $\mu=E[X]$, and any $\delta>1$, $\Pr[X\geq(1+\delta)\mu]\leq e^{-\frac{\delta\mu}{2}}$, and any $0\leq\delta\leq 1$, $\Pr[X\geq(1+\delta)\mu]\leq e^{-\frac{\delta^{2}\mu}{3}}$ and $\Pr[X\leq(1-\delta)\mu]\leq e^{-\frac{\delta^{2}\mu}{2}}$.
Lemma 0.
If a transaction chooses $k$ objects randomly and uniformly (out of $s$ shards), then the maximum number of transactions accessing any object is $l\leq 2k+4\ln s$, with probability at least $1-1/s$.
Proof.
Consider an object $o_{z}\in{\mathcal{O}}$. Let $X^{(z)}_{i}$ be the random variable such that $X^{(z)}_{1}=1$ if transaction $T_{i}$ accesses object $o_{z}$. Let $Y_{z}=\sum_{j=i}^{s}X^{(z)}_{j}$, and let $\mu=E[Y_{z}]$ be the number of transactions that access object $o_{z}$.
Using the union bound, the probability that there exists an object $o_{z}$ such that $Y_{z}$ exceeds a certain value $c$ is at most the sum of the probabilities that each $Y_{z}$ exceeds $c$.
$\mathrm{Pr}(l\geq c)=\mathrm{Pr}(\text{exists $o_{z}$ s.t. }Y_{z}\geq c)\leq\sum_{z=1}^{s}\mathrm{Pr}(Y_{z}\geq c)\ .$
Since each transaction accesses $k$ objects, and each shard has a transaction, the expected number (average) of transactions accessing each object is $\frac{s\cdot k}{s}=k$.
From Lemma 1 (Chernoff bound) for a random variable $Y_{z}$ and any $\delta>1$ with expected value $\mu$ is:
$\mathrm{Pr}(Y_{z}\geq(1+\delta)\mu)\leq e^{-\frac{\delta\mu}{2}}\ .$
In our case, $Y_{z}$ represents the number of transactions accessing a specific object, and $\mu=k$ (the expected number of transactions accessing each object).
We want to find the probability that the number of transactions accessing this object is greater or equal to $c$, i.e., $\mathrm{Pr}(Y_{z}\geq c)$. To apply the Chernoff bound, we set $c=(1+\delta)k$, which gives us $\delta=\frac{c-k}{k}$ and $\mu=k$. Let’s substitute these into the Chernoff bound:
$\mathrm{Pr}(Y_{z}\geq c)\leq e^{-\frac{\frac{(c-k)}{k}}{2}k}=e^{-\frac{(c-k)}{2}}\ .$
For the union bound, we sum these probabilities over all objects:
$\mathrm{Pr}(l\geq c)\leq\sum_{i=1}^{s}\mathrm{Pr}(Y_{z}\geq c)=s\cdot e^{-\frac{(c-k)}{2}}\ .$
We want this probability to be less than or equal to $\frac{1}{s}$, so:
$s\cdot e^{-\frac{(c-k)}{2}}\leq\frac{1}{s}\ .$
Solving for $c$, we get:
$c\geq k+4\ln s\ .$
We assume $c=(1+\delta)k$, where $\delta>1$. Thus, we can take
$c\geq 2k+4\ln s\ .$
Therefore, $l\leq 2k+4\ln s$ with probability at least $1-\frac{1}{s}$.
Lemma 0 (Lower Bound).
The time to process the transactions in bucket $B_{i}$ is at least $2^{i}$.
Proof.
This follows immediately from the definition of bucket $B_{i}$, since there is a transaction in $B_{i}$ that has distance at least $2^{i}$ to one of its accessing objects. ∎
Lemma 0 (Upper Bound).
If each transaction accesses at most $k$ objects randomly from different shards, then the time to process the transactions in any bucket $B_{i}$ is at most $O(2^{i}k\log D\cdot(k+\log s))$.
Proof.
Suppose $|B_{i}|$ is the total number of transactions in the bucket $B_{i}$. Let ${\mathcal{T}}_{x}^{b_{i}}$ denote the set of transactions in $B_{i}$ that use object $o_{j}\in{\mathcal{O}}$. Let $l_{x}=|{\mathcal{T}}_{x}^{b_{i}}|$ and $l_{i}=max_{x}l_{x}$ be the maximum number of transactions that used any object in ${\mathcal{O}}$. From Lemma 2, if the transaction accesses the objects randomly, the maximum degree of any object is bounded by $2k+4\ln s$. Thus, we can write $l_{i}\leq 2k+4\ln s$. Since all the transactions in the bucket $B_{i}$ access at most $k$ objects and each object is used by at most $l_{i}$ transactions. Thus, the maximum degree in the transaction graph of bucket $B_{i}$ is at most $kl_{i}$, and therefore, the conflict graph $G_{{\mathcal{T}}}$ can be greedily colored with $\xi\leq kl_{i}+1$ colors.
The distance between a transaction and its accessing object of bucket $B_{i}$ is at most $2^{i+1}-1<2^{i+1}$. Moreover, the simulation of Algorithm 1 has at most $4$ rounds of interactions between the transaction and its accessing objects for the confirmation and commit. Therefore, each color corresponds to $4\cdot 2^{i+1}$ rounds. Thus, it takes $4\cdot 2^{i+1}(kl_{i}+1)=O(2^{i}kl_{i})=O(2^{i}k(k+\log s))$ rounds to confirm and commit transactions.
Thus, if we have only one bucket $B_{i}$, then the time to process all the transactions of the bucket $B_{i}$ is at most:
(1) Maximum processing time for bucket $\text{Maximum processing time for bucket $B_{i}$}=2^{i+3}(kl_{i}+1)$
In our algorithm, transactions from lower index buckets are prioritized for processing over higher index buckets. As we have at most $\lfloor\log D\rfloor+1$ buckets, we analyzed the effect of it to process all the transactions.
Assume bucket $B_{i}$ is the worst bucket, and it takes more time to process all the transactions on its bucket than any other bucket. From equation 1, if we only have bucket $B_{i}$, the cost required for the scheduler to process all the transactions is at most $B_{i}^{c}=4\cdot 2^{i+1}(kl_{i}+1)$ rounds. However, if we consider all the buckets, the scheduler needs to process transactions from bucket 00 to bucket $\log D$. Thus, the total number of rounds needed to process all bucket transactions is at most:
(2) $\sum_{x=o}^{\log D}B_{x}^{c}\leq 2B_{i}^{c}\log D=2\cdot 4\cdot 2^{i+1}(kl_{i}+1)\log D=2^{i+4}(kl_{i}+1)\log D$
Thus, the total processing time to process all the transactions from all the buckets is at most $2^{i+4}(kl_{i}+1)\log D=O(2^{i}kl_{i}\log D)=O(2^{i}k\log D\cdot(k+\log s))$. ∎
From Equation 1, if we consider only one bucket $B_{i}$ and $B_{i}$ is the worst bucket, then total transactions processing time is at most $2^{i+3}(kl_{i}+1)$. However, from Equation 2, if we consider all the buckets, then total transactions processing time is at most $2^{i+4}(kl_{i}+1)\log D$. Thus, if we use a bucketing approach, then we need to pay an extra $O(\log D)$ cost over the worst bucket. Hence, we have the following corollary.
Corollary 0.
If we use a bucketing approach, we need to pay an extra $O(\log D)$ cost over the worst bucket.
Theorem 6.
For the General graph, if the transaction accesses $k$ objects randomly from different shards, then Algorithm 2 gives an $O(k\log D\cdot(k+\log s))$ approximation to the optimal schedule with probability at least $1-1/s$.
Proof.
From Lemma 4, the time to process all the transactions in the bucket $B_{i}$ is at most $O(2^{i}k\log D\cdot(k+\log s))$. Moreover, from Lemma 3, $2^{i}$ is the lower bound, which gives us $O(k\log D\cdot(k+\log s))$ approximation to the optimal schedule. ∎
5.2. Analysis for Line Graph with Buckets
We consider a graph as a line graph with $(V,E)$ where shards are connected in a line. We consider the sequence of $s$ shards, where $V=\{v_{1},v_{2},v_{3},\cdots,v_{s}\}$, and there is a edge between two vertices $v_{i},v_{i+1}$ of weight $1$ for $11\leq i\leq s-$ 1. We consider the position of the shards in the line graph to be left to right. Thus, $v_{1}$ is the leftmost shard, and $v_{s}$ is the rightmost shard. We can also consider mapping the line graph into the clique graph we discussed in the previous section. That means we can map every shard of the line to the corresponding shard in a clique graph, and every edge of a line graph corresponds to some edge in a clique graph, which connects direct neighbors in the line graph. Additional edges in the clique graph that connect indirect neighbors in the line graph have weights equal to the distance between the neighbors.
Lemma 0 (Lower Bound).
For the Line graph, the time to process the transactions in bucket $B_{i}$ is at least $2^{i}+\frac{l_{i}^{2}}{4}$, where $l_{i}$ is the maximum number of transactions in $B_{i}$ that access any object $o_{j}\in{\mathcal{O}}$.
Proof.
Suppose shards are arranged in a line, and each shard contains one transaction to be processed. Let $|B_{i}|$ be the total number of transactions in the bucket $B_{i}$. Suppose there is an object $o_{j}\in{\mathcal{O}}$, which is accessed by all the transactions that are in $B_{i}^{\prime}\subseteq B_{i}$. Thus the maximum number of transactions $l_{i}$ in $B_{i}$ that access an object $o_{j}$ is $|B_{i}^{\prime}|$, i.e. $l_{i}=|B_{i}^{\prime}|$.
We can decompose the bucket $B_{i}^{\prime}$ into partial buckets $\bar{B}^{\prime}_{0,i},\bar{B}^{\prime}_{1,i},\cdots,\bar{B}^{\prime}_{z,i}$ where partial bucket $\bar{B}^{\prime}_{z,i}$ store the transaction which access object $o_{j}$ from distance $[2^{z},2^{z+1})$.
Without loss of generality, suppose the object $o_{j}$ is in the middle of the line. Since each shard hosts a single transaction, the size of the bucket is bounded by the distance that comprised transactions have to traverse $|B_{z,i}^{\prime}|<2^{i+1}$. Thus $z\leq\log l_{i}$. Now calculate the number of transactions that are at a distance of at least $z-2\Longrightarrow\log l_{i}-2$.
$\sum_{z=0}^{\log l_{i}-2}|B_{z,i}^{\prime}|<\sum_{z=0}^{\log l_{i}-2}2^{z}\leq 2\cdot 2^{\log l_{i}-2}\leq l_{i}/2\ .$
As the transactions are in line, the remaining $l_{i}/2$ transactions must be in the bucket index $\bar{B}^{\prime}_{>z-2,i}$. Consider a bucket $\bar{B}^{\prime}_{z-1,i}$ which stores the transaction accessing object $o_{j}$ from distance $22^{z-1}\Longrightarrow 2^{\log l_{i}-1}=l_{i}/$ 2. This implies that other $l_{i}/2$ transactions are at least $l_{i}/2$ distance from object $o_{j}$. Thus, the maximum degree for the object $o_{z}$ is at least $l_{i}/2\cdot l_{i}/2=\frac{l_{i}^{2}}{4}$ because each $l_{i}/2$ conflicting transaction needs to be serialized, and they are far from object $o_{j}$ and need to be travel at least $l_{i}/2$ distance.
Moreover, there is at least one transaction in bucket $B_{i}$, which accesses the object at a distance of at least $2^{i}$; that is why they are in bucket $B_{i}$. Consequently, the total time to process all the transactions of bucket $B_{i}$ is at least $2^{i}+\frac{l_{i}^{2}}{4}$. ∎
Lemma 0 (Upper Bound).
For the Line graph, the time to process the transactions in bucket $B_{i}$ is at most $O(2^{i}kl_{i}\log D)$, where $l_{i}$ is the maximum number of transactions in $B_{i}$ that access any object $o_{j}\in{\mathcal{O}}$.
Proof.
Suppose shards are arranged in a line and consist of transactions that need to be processed in each shard. Let $|B_{i}|$ be the total number of transactions in the bucket $B_{i}$. In the worst case, all of the transactions can access the same object $o_{j}$. Thus, the maximum degree of the object $o_{j}$ access by transactions in bucket $B_{i}$ is $l_{i}=|B_{i}|$. Since all the transactions in the bucket $B_{i}$ access at most $k$ objects and each object is used by at most $l_{i}$ transactions. The maximum degree of the transaction graph of bucket $B_{i}$ is at most $kl_{i}$, and therefore, the transaction graph can be greedily colored with $\xi\leq kl_{i}+1$ colors.
The distance between a transaction and its accessing objects of bucket $B_{i}$ is at most $2^{i+1}-1<2^{i+1}$. Moreover, for each color, Algorithm 1 takes $4\cdot 2^{i+1}$ rounds to confirm and commit the transactions. Thus, transactions of the bucket $B_{i}$ take at most $4\cdot 2^{i+1}(kl_{i}+1)=O(2^{i}kl_{i})$ rounds to process transactions.
Additionally, from Corollary 5, if we consider all the buckets, we need to pay an extra $O(\log D)$ cost. Thus, the total execution time to process all the transactions of the bucket $B_{i}$ is at most $O(2^{i}kl_{i}\log D)$. ∎
Theorem 9 (Approximation).
If the shards are in a line graph and each transaction accesses $k$ objects, then Algorithm 2 provides an $O(k\sqrt{d}\log D)$ approximation to the optimal schedule, where $d$ is the worst distance of transaction and its accessing objects and $D$ diameter of the line graph.
Proof.
From Lemma 8, time to process all the transactions in any bucket $B_{i}$ is at most $O(2^{i}kl_{i}\log D)$. Moreover, from Lemma 7, lower bound to process all the transaction of bucket $B_{i}$ is at least $2^{i}+l_{i}^{2}/4=O(2^{i}+l_{i}^{2})$, which gives us $O(\frac{2^{i}kl_{i}\log D}{2^{i}+l_{i}^{2}})$ approximation. we will analyze two cases:
(i) When $l_{i}<2^{i/2}$:
$O(\frac{2^{i}kl_{i}\log D}{2^{i}+l_{i}^{2}})\leq O(\frac{2^{i}k2^{i/2}\log D}{2^{i}+{2^{i/2}}^{2}})<O(\frac{2^{i}k2^{i/2}\log D}{2^{i}})=O(k\sqrt{2^{i}}\log D)\ .$
We can replace $2^{i}$ with $d$ because $2^{i}$ is the distance between the home shard of the transactions and its accessing objects, and $d$ is the maximum distance between transactions and their accessing objects in the line graph (L). Thus, we will get $O(k\sqrt{d}\log D)$.
(ii) When $l_{i}\geq 2^{i/2}$:
$O(\frac{2^{i}kl_{i}\log D}{2^{i}+l_{i}^{2}})\leq O(\frac{2^{i}kl_{i}\log D}{l_{i}^{2}})=O(\frac{2^{i}k\log D}{l_{i}})=O(k\sqrt{2^{i}}\log D)$
$=O(k\sqrt{d}\log D)\ .$
From the above analysis and expression, we can say that the line graph Algorithm 1 gives an $O(k\sqrt{d}\log D)$ approximation to optimal schedule. ∎
6. Distributed Scheduler (DS)
The scheduling algorithm we presented earlier uses a central authority in a shard with knowledge about all current transactions and the maximum degree of the transaction graph. However, such a central authority might not exist in blockchain sharding since each transaction is generated in a distributed manner in any shard. Here, we discuss a fully distributed scheduling approach using a clustering technique that allows the transaction schedule to be computed in a decentralized manner without requiring a central authority in the shard. We can also view distributed scheduling Algorithm 3 as a simulation of centralized scheduling Algorithm 1, which is described in Subsection 6.3.
6.1. Shard Clustering
In the distributed scheduler, shards are distributed across the network, and the distance between the home shard of the transaction and its accessing objects (destination shards) ranges from $1$ to $d\leq D$, where $D$ is the diameter of the shard graph. Let us suppose shards graph $G_{s}$ constructed with $s$ shards, where the weights of edges between shards denote the distances between them. We consider that $G_{s}$ is known to all the shards. We define $z$-neighborhood of shard $S_{i}$ as the set of shards within a distance of at most $z$ from $S_{i}$. Moreover, the 0-neighborhood of shard $S_{i}$ is the $S_{i}$ itself.
We consider that our distributed scheduling algorithm uses a hierarchical decomposition of $G_{s}$ which is known to all the shards and calculated before the algorithm starts. This shard clustering (graph decomposition) is based on the clustering techniques in (Gupta et al., 2006) and which were later used in (Sharma and Busch, 2014; Busch et al., 2022; Adhikari et al., 2024a, b). We divide the shard graph $G_{s}$ into the hierarchy of clusters with $H_{1}=\lceil\log D\rceil+1$ layers (logarithms are in base 2), and a layer is a set of clusters, and a cluster is a set of shards. Layer $q$, where $0\leq q<H_{1}$, is a sparse cover of $G_{s}$ such that:
  • Every cluster of layer $q$ has (strong) diameter of at most $O(2^{q}\log s)$.
  • Every shard participates in no more than $O(\log s)$ different clusters at layer $q$.
  • For each shard $S_{i}$ there exists a cluster at layer $q$ which contains $(2^{q}-1)$- neighborhood of $S_{i}$ within that cluster.
The clusters have a strong diameter, which measures distances within a cluster on the induced subgraph of $G_{s}$.
For each layer $q$, the sparse cover construction in (Gupta et al., 2006) is actually obtained as a collection of $H_{2}=O(\log s)$ partitions of $G_{s}$. These $H_{2}$ partitions are ordered as sub-layers of layer $q$ labeled from 00 to $H_{2}-1$. A shard might participate in all $H_{2}$ sub-layers but potentially belongs to a different cluster at each sub-layer. At least one of these $H_{2}$ clusters at layer $q$ contains the whole $12^{q}-$ 1 neighborhood of $S_{i}$.
In each cluster at layer $q$, a leader shard is specifically designated such that the leader’s $(2^{q}-1)$-neighborhood is in that cluster. As we give an idea of layers and sub-layers, we define the concept of height as a tuple $h=(h_{1},h_{2})$, where $h_{1}$ denotes the layer and $h_{2}$ denotes the sub-layer. Similar to (Sharma and Busch, 2014; Busch et al., 2022; Adhikari et al., 2024a, b), heights follow lexicographic order.
The home cluster for each transaction $T_{i}$ is defined as follows: suppose $S_{i}$ is the home shard of $T_{i}$, and $z$ is the maximum distance from $S_{i}$ to the destination shards that will be accessed by $T_{i}$; the home cluster of $T_{i}$ is the lowest-layer (and lowest sub-layer) cluster in the hierarchy that contains $z$-neighborhood of $S_{i}$. Each home cluster consists of one dedicated leader shard, which will handle all the transactions that have their home shard in that cluster (i.e., transaction information will be sent from the home shard to the cluster leader shard to determine the schedule).
6.2. Distributed Scheduling Algorithm
In the distributed scheduler Algorithm 3, each home shard determines the home cluster for its transaction based on the distance between the home shard of the transaction and its accessing objects (destination shards) and sends that transaction information to the leader of the respective home cluster. Each cluster $C_{z}$ belongs to some layer $q$ and sublayer $r$, where $0\leq q<H_{1}$ and $0\leq r<H_{2}$; Suppose a transaction $T_{i}$ with home cluster $C_{z}$; we also say that $T_{i}$ is at level $(q,r)$. The leader shard of $C_{z}$ assigns an integer color to each transaction $T_{i}$. We define the concept of the height for $T_{i}$ represented by a tuple $(q,r,\text{color})$. The heights of transactions are ordered lexicographically, and a the priority scheme is implemented. The priority for committing transactions is determined based on this order, giving precedence to transactions generated in lower cluster layers. The reason behind this priority scheme is to take the benefit of the locality to improve the transaction latency. Note that a home shard may have transactions at various levels. Those transactions will be processed at the respective level.
For each level $(q,r)$, each cluster $C_{z}$ invokes the Algorithm 1 parallelly to schedule and process the transactions. While invoking Algorithm 1 for each cluster $C_{z}$ we set the parameters as $G[C_{z}]$, ${\mathcal{T}}(C_{z})$, $\ell(C_{z})$, where $G[C_{z}]$ is the cluster graph induced by $C_{z}$, ${\mathcal{T}}(C_{z})$ is the set of transactions whose home cluster is $C_{z}$, and $\ell(C_{z})$ is the leader shard of the cluster $C_{z}$ which is also presented in Algorithm 3.
Input : Shard graph $G_{s}$; set of txns ${\mathcal{T}}$;
1
2Assume a hierarchical cluster decomposition of $G_{s}$, which is known to all the shards in $G_{s}$;
3
4Each cluster $C_{z}$ belongs at some level $(q,r)$, at layer $q$ and sublayer $r$ of the hierarchical decomposition, where $0\leq q<H_{1}$ and $0\leq r<H_{2}$; the levels $(q,r)$ are ordered lexicographically;
5
6The home shard of each txn $T_{i}\in{\mathcal{T}}$ discovers the position of its destination shards in $G_{s}$;
7
8The home shard of each txn $T_{i}$ picks as a home cluster as the lowest level $(q,r)$ cluster $C_{z}$, which includes $T_{i}$ and the destination shards of its subtransactions;
9
10
11/* Run Algorithm 1 for each cluster */
12
13 for each cluster layer $q\leftarrow 0$ to $H_{1}$ and sublayer $r\leftarrow 0$ to $H_{2}$ do
14       for each cluster $C_{z}$ in level $(i,j)$ in parallel do
15             ${\mathcal{T}}(C_{z})\leftarrow$ set of transactions with home cluster $C_{z}$;
16             $\ell(C_{z})$leftarrow leader shard of cluster $C_{z}$;
17             $G[C_{z}]\leftarrow$ cluster graph induced by $C_{z}$;
18             Run CentralizedScheduler($G[C_{z}]$, ${\mathcal{T}}(C_{z})$, $\ell(C_{z})$);
19
20
21
Algorithm 3 DistributedScheduler
6.3. Simulating Centralized Scheduler
We can use the centralized scheduler algorithm (Algorithm 1) in each cluster of the distributed scheduler. In Algorithm 1, each leader shard has knowledge of transactions assigned to it and schedules them in a conflict-free manner using coloring. Similarly, in the distributed scheduler, each cluster leader shard knows all the transactions within its cluster and determines the transaction schedule by coloring the transactions conflict graph.
For example, let’s suppose the shard cluster graph has a maximum diameter $D$. Assume the distance between all the transactions and their accessed objects is $d\leq D$ far away from each other. In this scenario, each home shard of the transaction reports to the same home cluster, and the leader shard of that home cluster determines the schedule by coloring all the transactions similar to Algorithm 1. However, the distributed scheduler algorithm can provide significant benefits if transactions access nearby objects, allowing for parallel processing at each cluster level. Moreover, the distributed scheduler algorithm is based on hierarchical clustering, which needs to pay some constant factor delay due to layer and sublayer hierarchy with a priority scheme.
Moreover, to improve the bounds for the Line graph and for special cases in the general graph when the $k$ objects are selected randomly in the distributed setting, we can simulate the Centralized Scheduler with Bucket (Algorithm 2) in a distributed manner. For each cluster $C_{z}$ graph (induced graph), we can run Algorithm 2, which will ultimately simulate Algorithm 1.
Due to the layer and sublayer structure, we need to pay an $O(\log^{2}s)$ cost over the analysis obtained in the centralized approach, which is provided in the following Theorem 1.
Theorem 1.
The distributed scheduler has a competitive ratio of $O(\mathcal{A}_{\mathcal{CS}}\log^{2}s)$, where $\mathcal{A}_{\mathcal{CS}}$ is the approximation ratio of the centralized scheduling algorithm ($\mathcal{CS}$).
Proof.
If the transactions are scheduled by the centralized scheduling algorithm 1, then their execution time would be within the centralized scheduler from optimal. However, in the distributed approach, the schedule is restricted by the layer and sublayer hierarchy of the cluster, as the lowest layer gets priority over the higher layer.
Suppose we only have transactions with cluster layer $q_{i}$ having cluster diameter $d_{q}$. Let $S_{q}$ be the leader shard of cluster $q_{i}$. Let ${\mathcal{T}}_{w}$ denote the set of transactions in $S_{q}$ that use object $O_{w}$. Let’s denote $l_{w}=|{\mathcal{T}}_{w}|$ and $l=\max_{w}l_{w}$ Suppose each transaction uses $k$ objects, and every object is used by at most $l$ transactions. Thus, the maximum degree in the transactions graph is at most $kl$, and therefore, the transactions graph can be greedily colored with $\xi\leq kl+1$ colors. Moreover, the diameter of the cluster is $d_{q}$, and for each color, Algorithm 1 takes at most $4$ rounds of interactions between the home shard of transaction and its accessing objects (destination shards). Thus, it takes $4d_{q}(kl+1)=O(kld_{q})$ rounds to process the transactions. Since $\Omega(l)$ is the lower bound, it gives $O(kd_{q})$ approximation. Which is equal to the approximation of the centralized scheduler algorithm Theorem 3. Thus we can say $\mathcal{A}_{\mathcal{CS}}=O(kd)$ for diameter $d$.
Let $q^{\prime}$ be the maximum layer accessed by any transaction where the diameter is at most $d_{q}^{\prime}$. If we only have layer $q_{i}$, the scheduler will require $4d_{q}(kl+1)$ rounds. However, if we consider the distributed scheduler with the layer and sublayer hierarchy, $S_{q}$ needs to process transactions from all layers $0,1,\cdots,q^{\prime}$ and sublayers indexed by $r$, which goes from 00 to $H_{2}-1$.
As cluster layer $q$ has diameter $O(2^{q}\log s)$, which is discussed in the shard clustering subsection above. So $d_{q}=O(2^{q}\log s)$ and we can write $d_{q}=c_{1}2^{q}\log s$ for some constant $c_{1}$. This implies $\sum_{q=0}^{q^{\prime}}d_{q}\leq 2d_{q}$. Thus, the total number of rounds needed to schedule all level $(q,r)$ clusters at $S_{q}$ is at most
(3) $\sum_{q=o}^{q^{\prime}}\sum_{r=0}^{H_{2}-1}4d_{q}(kl+1)\leq 8d_{q}(kl+1)H_{2}\ .$
We can replace $H_{2}=c_{1}^{\prime}\log s$ and $d_{q}=c_{2}^{\prime}d\log s$. Let’s combine positive constant $c_{1}^{\prime}$ and $c_{2}^{\prime}$ as $c$. The equation 3 becomes $8\cdot cd(kl+1)\log^{2}s$. Since $\Omega(l)$ is the lower bound and $\mathcal{A}_{\mathcal{CS}}=O(kd)$, we conclude that the distributed scheduler has competitive ratio $O(\mathcal{A}_{\mathcal{CS}}\log^{2}s)$. ∎
From Theorem 1, we conclude that we need to pay an extra $O(\log^{2}s)$ rounds penalty for the distributed scheduler (Algorithm 3) due to hierarchical clustering. Hence, we have the following corollary:
  • For a General graph, Algorithm 3 provides an $O(kd\log^{2}s)$ approximation to the optimal schedule, where $d$ is the maximum distance between the transaction and its accessing objects.
  • Algorithm 3 provides an $O(k\log^{3}s)$ approximation for Hypercube, Butterfly, and $g$-dimensional grids graphs, where $g=O(\log s)$.
  • For a General graph, If the transaction accesses $k$ objects randomly, then Algorithm 3 gives an $O(k\log D\cdot(k+\log s)\log^{2}s)$ approximation to the optimal schedule with probability at least $1-1/s$, where $D$ is the diameter of shard graph.
  • For a Line graph, Algorithm 3 gives an $O(k\sqrt{d}\log D\cdot\log^{2}s)$ approximation to the optimal schedule, where $k$ is the maximum number of objects accessed by a transaction.
7. Conclusions
We presented efficient and fast execution time schedules to process batches of transactions for blockchain sharding systems under a synchronous communication model. Our approach includes centralized and distributed schedulers. In the centralized scheduler, we assume each shard has global transaction information, while in the distributed scheduling algorithm, shards do not have global information. We provide optimal approximations for the proposed scheduling algorithm, along with lower and upper bounds. Our results represent the first known attempts to obtain provably fast batch transaction scheduling for blockchain sharding.
For future work, it would be interesting to extend our results to the online and dynamic setting, where the set of transactions to be processed is continuously generated and injected into a shard. Additionally, exploring efficient communication mechanisms between shards and considering the impact of network congestion, where network links have bounded capacity, would be interesting areas for further research.
Acknowledgements.
This paper is supported by NSF grant CNS-2131538.